OpenStudy (anonymous):

The force that should be applied on a 5kg body so that it has a downward acceleration of 4m/s^2?

OpenStudy (anonymous):

Force=mass*acceleration=5*4=20N

OpenStudy (anonymous):

The options are 29N upward 69Ndownwards...

OpenStudy (apoorvk):

in general it has an acc. of 10m/sec^2 under free-fall. then the force acting would be mg. Now an opposing force acts such, that it reduces the net force to '4m' downwards, where '4' is its acceleration. So, what do you think now?

OpenStudy (anonymous):

@apoorvk is the ans 29 upwards?

OpenStudy (apoorvk):

if g=9.8, yeah 29N upwards.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but how??

OpenStudy (anonymous):

mass*net acceleraton=m*10-force 5*4=5*9.8-F F=m*(9.8-4)=29N =2

OpenStudy (anonymous):

FORGET ABOUT THE LAST LINE.IT'S MY MISTAKE!!!

OpenStudy (anonymous):

Thanzzz